# Symbolic Planning and Multi-Agent Path Finding in Extremely Dense Environments with Movable Obstacles

### v1

@@ Abstract
We introduce the Block Rearrangement Problem (BRaP), a challenging component of large warehouse management which involves rearranging storage blocks within dense grids to achieve a target state. We formally define the BRaP as a graph search problem. Building on intuitions from sliding puzzle problems, we propose five search-based solution algorithms, leveraging joint configuration space search, classical planning, multi-agent pathfinding, and expert heuristics. We evaluate the five approaches empirically for plan quality and scalability. Despite the exponential relation between search space size and block number, our methods demonstrate efficiency in creating rearrangement plans for deeply buried blocks in up to $\times80$ grids.
@@ 1 Introduction
Amazon Robotics fulfillment centers are large warehouses that utilize robotic Automated Storage and Retrieval Systems to store inventory and fulfill customer orders. As illustrated in Figure 1, fulfillment centers store inventory items in shelves (blocks) arranged into storage grids separated by travel lanes. Robot drives can lift and move these blocks. When a station at the periphery requests a target block for item picking or stowing, robot drives move that block, transport it to the storage grid’s boundary and then utilize travel lanes to bring it to the station. The system leverages dense storage grids to achieve efficient space utilization and lower fulfillment costs. This creates a challenge: target blocks may be deeply buried and obstructed by other blocks, requiring complex rearrangement plans.
@@
Figure 1: Fulfillment center with blocks in storage grids separated by travel lanes. Target blocks marked in orange.
@@
We define the Block Rearrangement Problem (BRaP) as follows. Given a grid layout containing target blocks, non-target blocks, empty cells, and obstacles, as illustrated in Figure 2a, the objective is to find the lowest-cost plan to move each target block to a set of desired goal locations. Note that the non-target blocks are also referred to as movable obstacles in existing work (Papadimitriou et al. 1994). These goal locations can be either contiguous or disjoint. When goal locations are at the grid boundary, target blocks are rearranged for direct travel lane access. Alternatively, when goal locations are within the grid, target blocks are positioned for potential future access. The plan consists of block moving actions with their associated timestamps. Cost metrics can include sum of action costs, makespan, or other user-defined objectives. An example plan for the problem shown in Figure 2a is illustrated in Figure 2b.
The Block Rearrangement Problem in dense warehouses shares fundamental structural similarities with various puzzles and real-world problems, including the sliding tile puzzle (Gozon and Yu 2024), Rush Hour puzzle (Cian et al. 2022), block relocation problem (Lu, Zeng, and Liu 2020), box world/Sokoban (Zawalski et al. 2024), parking lot rearrangement (Guo and Yu 2023), and Multi-Agent Path Finding (MAPF) (Stern et al. 2019; Li et al. 2022; Shen et al. 2023; Okumura 2023b). These problems involve rearrangement in discrete space, aiming to achieve goal configurations under movement constraints. Due to the combinatorial increase in the number of possible configurations, these problems are generally NP-hard (Ratner and Warmuth 1986; Flake and Baum 2002; Yu and LaValle 2013a).
@@ 2 Problem Description
A BRaP is defined on a graph $G=(V,E)$ with the set of vertices $V$ and edges $E=\{(u,v)\ |\ u,v\in V\}$. Let $\mathcal{I}=\{1,\cdots,n_{\mathcal{I}}\}$ denote the set of target blocks (marked in orange in Figure 2) and each target $i\in\mathcal{I}$ starts at an initial vertex $s_{i}\in V$ and must be moved to a set of desired goal vertices $V_{i}\subseteq V$. Additionally, $\mathcal{J}=\{1,\cdots,n_{\mathcal{J}}\}$ denotes non-target blocks without assigned goals. Actions $A$ comprise a set of operations that modifies the time or location of the blocks. A path $p_{i}$ for block $i\in\mathcal{I}\cup\mathcal{J}$ consists of an action sequence $(a_{i}^{1},a_{i}^{2},\cdots,a_{i}^{|p_{i}|}),a_{i}^{k}\in A$. Target block paths must move targets to goal vertices to be completed, while non-target block paths facilitate target block movements. A solution $P=\{p_{i}\ |\ i\in\mathcal{I}\cup\mathcal{J}\}$ to a BRaP is a set of conflict-free paths, that transition all the target blocks to their goal vertices. The objective is to find a feasible solution minimizing either sum of action costs, makespan, or other user-defined cost metrics, as detailed in Sec. 2.3.
@@ 2.1 Actions
It is important to note that, according to the above definition, the destination of a move action must be an empty space. We prohibit platooning, where one block closely follows another, to better align with warehouse operational constraints. Specifically, the movement must avoid the following three conflicts (Stern et al. 2019):
@@ 2.2 Terminal State
The objective is to empty the target block set through completion actions. Three distinct completion actions are possible for target blocks. In the first case, a target block reaching the goal (e.g., the boundary) is considered rearranged and removed, converting its location to free space. In the second case, a target block becomes a non-target block upon reaching its goal location. In the third case, a target block becomes an immovable obstacle at its goal location. Each completion action corresponds to a specific practical scenario. This paper focuses on developing algorithms and conducting experiments using the third option.
@@ 2.3 Objectives and Metrics
Multiple metrics can be used to evaluate the quality of the BRaP solutions, including the number of moves of target blocks, the number of moves of non-target blocks, sum of action costs, and the makespan of the whole plan. In Block Rearrangement, we care about all of these metrics.
@@
Block type move wait complete
Target 2 1 2
Non-target 2 0 N/A
@@
We calculate the objective metrics: composite cost of a solution $P=\{p_{i}\ |\ i\in\mathcal{I}\cup\mathcal{J}\}$ and the makespan of the solution, according to (2)-(3), respectively.
@@ 3 Related Work
As mentioned in Sec. 1, BRaP is closely related to various puzzles and research problems. Solutions to these problems have employed diverse approaches including symbolic planning (Davesa Sureda et al. 2024), search and conflict resolution (Li et al. 2022; Sharon et al. 2015; Ma et al. 2019), mixed-integer programming (Guo and Yu 2023; Lu, Zeng, and Liu 2020), boolean satisfiability (Cian et al. 2022), reinforcement learning (Shoham and Elidan 2021; Damani et al. 2021; Wang et al. 2025), and supervised learning formulations (Li et al. 2020; Jiang et al. 2025) to find low-cost plans. Given their demonstrated optimality and efficiency, we leverage symbolic planning and MAPF algorithms to develop the planning methods in this paper.
@@ 3.1 Symbolic Planning and PDDL
Symbolic planning is a model-based approach in artificial intelligence that solves decision-making problems by explicitly representing states, actions, and their transitions using symbolic logic. Planning languages like Planning Domain Definition Language (PDDL) enable the formulation of complex planning problems as structured domains, encompassing crucial elements such as initial conditions, feasible actions, and desired goal states, allowing planners to search for action sequences that achieve specified goals. PDDL has been successfully applied to various domains: modeling Rush Hour and sliding puzzles (Davesa Sureda et al. 2024), formulating Rubik’s cube solutions (Muppasani, Pallagani, and Srivastava 2024), solving logistics problems for transport routing (Helmert 2009), and finding robot paths in grid worlds (Estivill-Castro and Ferrer-Mestres 2013). Symbolic representations have also proven effective in task and motion planning for robotic manipulation (Garrett et al. 2021; Silver et al. 2021). Recent work has combined the symbolic representation with deep learning and large language models to create flexible long-horizon plans (Srivastava et al. 2022; Agia et al. 2023; Lin et al. 2023). Building on this foundation, we apply symbolic planning in configuration space and PDDL descriptions to develop planning algorithms for BRaPs.
@@ 3.2 Multi-Agent Path Finding
Multi-Agent Path Finding (Stern et al. 2019) is a problem of navigating a team of agents from their start locations to goal locations without collisions. The BRaP can be modeled as a complex MAPF variant that combines elements from the classic MAPF problem (Stern et al. 2019), where agents must be moved to dedicated goal locations, the anonymous MAPF problem (Stern et al. 2019; Yu and LaValle 2013b), where goal vertices are unlabeled and are interchangeable for target blocks, and the Graph Motion Planning with Movable Obstacles problem (Papadimitriou et al. 1994), where numerous non-target blocks act as dynamic obstacles that must be moved to clear paths. But different from the anonymous MAPF problem, target blocks may been assigned distinct goal vertex sets thus goal vertices are not fully interchangeable. The BRaP is made further challenging by the high-density environment, where non-target blocks often occupy the vast majority of the workspace, and by its specific operational constraints prohibiting following. Although following conflicts are defined in the literature, most solvers ignore them and it is non-trivial to incorporate them into certain MAPF algorithms. These combined factors mean that even though powerful, state-of-the-art solvers like LaCAM (Okumura 2023b, a, 2024) are effective for high-density MAPF problems, they are not directly applicable to BRaPs. Therefore, building upon these advanced frameworks, we develop a MAPF-based solution tailored to the specific complexities of the BRaP.
@@ 4.1 Configuration Space Search
The BRaP can be formulated as a graph search in configuration space. Starting from an initial grid configuration, available actions transform the state until reaching a goal configuration that satisfies desired criteria. Actions at each time step represent edges that transition between neighboring configuration states. Each state comprises time, target set (vertices containing target blocks), free space set (unoccupied vertices), and completed target set (vertices containing completed target blocks).
An example graph of configuration states and actions is illustrated in Figure 4. According to the above definition, the root state in the graph is: time: 0, targets: {(2,1), (2,3), (2,6)}, free space: {(3,1), (1,2), (2,2), (2,5)}, completed target: $\emptyset$. The goal state should contain an empty target set.
@@
The configuration space can be modified using the three actions described in Sec. 2.1. To ensure computational tractability and reduce the branching factor, we initially restrict the system to one action per time step. This restriction renders the wait action redundant and thus disallowed. Methods supporting multiple actions at each time step will be discussed in subsequent sections.
Based on the definition of the states and actions, a configuration graph can be constructed. An example is illustrated in Figure 4. With the cost of the actions defined, graph search techniques like Dijkstra and A* algorithms can be applied to find the minimum cost path from the initial state to the goal state (where there is no longer any target blocks).
A* with an admissible heuristic function is guaranteed to be complete (always find a solution if there is) and optimal. We formulate an admissible heuristic by assuming target blocks move along least-blocking vertices, with each blocking block requiring only one move action for removal. The detailed heuristic formulation is presented in Appendix B.
@@ 4.2 Priority-based Configuration Space Search
In contrast to the single action per time step approach of Sec. 4.1, we now develop a formulation allowing one action per target block per time step, enabling parallel execution and reducing rearrangement makespan.
Algorithm 1 begins by assigning priorities based on block proximity to goals, as our experiments showed this produces better solutions. The priority for each target $i\in\mathcal{I}$ is determined using the heuristic $h(s,i)$ described in Appendix B, where smaller values indicate higher priorities.
The search iteratively creates rearrangement plans for individual targets while respecting plans of higher-priority blocks as constraints. For a prior plan $P$, the computeConstraint($P$) function generates constraints ensuring the preconditions of higher-priority actions are satisfied. For example, if move(i,t,u,v)(i,t,u,v) appears in $P$, a constraint in $\xi$ ensures block $i$ occupies $u$ at time $t$ while $v$ remains unoccupied, and no other actions can modify these vertices at time $t$. The computeConstrainedPlan() function generates plans by pruning states violating $\xi$ and treating prior plans as external forces, ensuring parallel executability of the resulting plans.
@@
2Input: Grid $G(V,E)$, target set $\mathcal{I}$, goal vertices $V_{i}(i\in\mathcal{I})$

### v2

# Symbolic Planning and Multi-Agent Path Finding in Extremely Dense Environments with Unassigned Agents

@@ Abstract
We introduce the Block Rearrangement Problem (BRaP), a challenging component of large warehouse management which involves rearranging storage blocks within dense grids to achieve a goal state. We formally define the BRaP as a graph search problem. Building on intuitions from sliding puzzle problems, we propose five search-based solution algorithms, leveraging joint configuration space search, classical planning, multi-agent pathfinding, and expert heuristics. We evaluate the five approaches empirically for plan quality and scalability. Despite the exponential relation between search space size and block number, our methods demonstrate efficiency in creating rearrangement plans for deeply buried blocks in up to $\times80$ grids.
@@ 1 Introduction
Amazon Robotics fulfillment centers are large warehouses that utilize robotic Automated Storage and Retrieval Systems to store inventory and fulfill customer orders. As illustrated in Figure 1, fulfillment centers store inventory items in shelves (blocks) arranged into storage grids separated by travel lanes. Robot drives can lift and move these blocks. When a station at the periphery requests an assigned block for item picking or stowing, robot drives move that block, transport it to the storage grid’s boundary and then utilize travel lanes to bring it to the station. The system leverages dense storage grids to achieve efficient space utilization and lower fulfillment costs. This creates a challenge: assigned blocks may be deeply buried and obstructed by other blocks, requiring complex rearrangement plans.
@@
Figure 1: Fulfillment center with blocks in storage grids separated by travel lanes. Assigned blocks marked in orange.
@@
We define the Block Rearrangement Problem (BRaP) as follows. Given a grid layout containing assigned blocks, unassigned blocks, empty cells, and obstacles, as illustrated in Figure 2a, the objective is to find the lowest-cost plan to move each assigned block to a set of desired goal locations. Note that the blocks are also referred to as agents in other problem settings (Stern et al. 2019). These goal locations can be either contiguous or disjoint. When goal locations are at the grid boundary, assigned blocks are rearranged for direct travel lane access. Alternatively, when goal locations are within the grid, assigned blocks are positioned for potential future access. The plan consists of block moving actions with their associated timestamps. Cost metrics can include sum of action costs, makespan, or other user-defined objectives. An example plan for the problem shown in Figure 2a is illustrated in Figure 2b.
The Block Rearrangement Problem in dense warehouses shares fundamental structural similarities with various puzzles and real-world problems, including the sliding tile puzzle (Gozon and Yu 2024), Rush Hour puzzle (Cian et al. 2022), block relocation problem (Lu et al. 2020), box world/Sokoban (Zawalski et al. 2024), parking lot rearrangement (Guo and Yu 2023), Multi-Agent Path Finding (MAPF) (Stern et al. 2019; Li et al. 2022; Shen et al. 2023; Okumura 2023b), and MAPF with Unassigned Agents (Felner and Stern 2026). These problems involve rearrangement in discrete space, aiming to achieve goal configurations under movement constraints. Due to the combinatorial increase in the number of possible configurations, these problems are generally NP-hard (Ratner and Warmuth 1986; Flake and Baum 2002; Yu and LaValle 2013b).
@@ 2 Problem Description
A BRaP is defined on a graph $G=(V,E)$ with the set of vertices $V$ and edges $E=\{(u,v)\ |\ u,v\in V\}$ connecting adjacent vertices. Let $\mathcal{I}=\{1,\cdots,n_{\mathcal{I}}\}$ denote the set of assigned blocks (marked in orange in Figure 2) and each assigned block $i\in\mathcal{I}$ starts at an initial vertex $s_{i}\in V$ and must be moved to a set of desired goal vertices $V_{i}\subseteq V$. Additionally, $\mathcal{J}=\{n_{\mathcal{I}}+1,\cdots,n_{\mathcal{I}}+n_{\mathcal{J}}\}$ denotes unassigned blocks without desired goals. Actions $A=\{\textit{move}{},\textit{wait}{},\textit{complete}{}\}$ comprise a set of operations that modifies the time or location of the blocks. A path $p_{i}$ for block $i\in\mathcal{I}\cup\mathcal{J}$ consists of an action sequence $(a_{i}^{1},a_{i}^{2},\cdots,a_{i}^{|p_{i}|}),a_{i}^{k}\in A$. Assigned block paths must move assigned blocks to goal vertices to be completed, while unassigned block paths facilitate assigned block movements. A solution $P=\{p_{i}\ |\ i\in\mathcal{I}\cup\mathcal{J}\}$ to a BRaP is a set of conflict-free paths, that transition all the assigned blocks to their goal vertices. The objective is to find a feasible solution minimizing either sum of action costs, makespan, or other user-defined cost metrics, as detailed in Sec. 2.3.
@@ 2.1 Actions
It is important to note that, according to the above definition, the destination of a move action must be an empty vertex. We prohibit platooning, where one block closely follows another, to better align with warehouse operational constraints. Specifically, the movement must avoid three conflicts (Stern et al. 2019):
@@ 2.2 Terminal State
The objective is to empty the assigned block set through completion actions. Three distinct completion actions are possible for assigned blocks. In the first case, an assigned block reaching the goal (e.g., the boundary) is considered rearranged and removed, converting its location to an empty vertex. In the second case, an assigned block becomes unassigned upon reaching its goal location. In the third case, an assigned block becomes an obstacle at its goal location. Each completion action corresponds to a specific practical scenario. This paper focuses on developing algorithms and conducting experiments using the third option.
@@ 2.3 Objectives and Metrics
Multiple metrics can be used to evaluate the quality of the BRaP solutions, including the number of moves of assigned blocks, the number of moves of unassigned blocks, sum of action costs, and the makespan of the whole plan. In Block Rearrangement, we care about all of these metrics.
@@
Block type move wait complete
Assigned 2 1 2
Unassigned 2 0 N/A
@@
We define the objectives: composite cost of a solution $P=\{p_{i}\ |\ i\in\mathcal{I}\cup\mathcal{J}\}$ and the makespan of the solution, according to (2)-(3), respectively.
@@ 3 Related Work
As mentioned in Sec. 1, BRaP is closely related to various puzzles and research problems. Solutions to these problems have employed diverse approaches including symbolic planning (Davesa Sureda et al. 2024), search and conflict resolution (Li et al. 2022; Sharon et al. 2015; Ma et al. 2019), mixed-integer programming (Guo and Yu 2023; Lu et al. 2020), boolean satisfiability (Cian et al. 2022), reinforcement learning (Shoham and Elidan 2021; Damani et al. 2021; Wang et al. 2025), and supervised learning formulations (Li et al. 2020; Jiang et al. 2025) to find low-cost plans. Given their demonstrated optimality and efficiency, we leverage symbolic planning and MAPF algorithms to develop the planning methods in this paper.
@@ 3.1 Symbolic Planning and PDDL
Symbolic planning is a model-based approach in artificial intelligence that solves decision-making problems by explicitly representing states, actions, and their transitions using symbolic logic. Planning languages like Planning Domain Definition Language (PDDL) enable the formulation of complex planning problems as structured domains, encompassing crucial elements such as initial conditions, feasible actions, and desired goal states, allowing planners to search for action sequences that achieve specified goals. PDDL has been successfully applied to various domains: modeling Rush Hour and sliding puzzles (Davesa Sureda et al. 2024), formulating Rubik’s cube solutions (Muppasani et al. 2024), solving logistics problems for transport routing (Helmert 2009), and finding robot paths in grid worlds (Estivill-Castro and Ferrer-Mestres 2013). Symbolic representations have also proven effective in task and motion planning for robotic manipulation (Garrett et al. 2021; Silver et al. 2021). Recent work has combined the symbolic representation with deep learning and large language models to create flexible long-horizon plans (Srivastava et al. 2022; Agia et al. 2023; Lin et al. 2023). Building on this foundation, we apply symbolic planning in configuration space and PDDL descriptions to develop planning algorithms for BRaPs.
@@ 3.2 Multi-Agent Path Finding
Multi-Agent Path Finding (Stern et al. 2019) is a problem of navigating a team of agents from their start locations to goal locations without collisions. The BRaP can be modeled as a complex MAPF variant that combines elements from the classic MAPF problem (Stern et al. 2019), where agents (blocks) must be moved to dedicated goal locations, the anonymous MAPF problem (Stern et al. 2019; Yu and LaValle 2013a), where goal vertices are unlabeled and are interchangeable for assigned blocks, and the Graph Motion Planning Problem (Papadimitriou et al. 1994), where unassigned blocks exists and must be moved to clear paths for assigned blocks but is limited to one motion per step. More recently, Felner and Stern (2026) introduced the Multi-Agent Path Finding with Unassigned Agents (MAPFUA), which formally defines this setting where assigned agents (our assigned blocks) have goals and others (our unassigned blocks) can move to clear the way. Our BRaP can be seen as a specific, challenging variant of MAPFUA, characterized by extremely dense environments, strict movement constraints, and task assignments. For a broader discussion of the MAPFUA problem and its potential variants, we direct readers to the MAPFUA paper (Felner and Stern 2026).
Different from the anonymous MAPF problem, assigned blocks may be assigned distinct goal vertex sets thus goal vertices are not fully interchangeable. This aspect is related to the Target Assignment and Path Finding (TAPF) problem (Ma and Koenig 2016), where agents in a team can be assigned to any of one of the team’s dedicated goal vertices. However, the TAPF problem does not include unassigned agents or blocks.
The BRaP is made further challenging by the high-density environment, where unassigned blocks often occupy the vast majority of the workspace, and by its specific operational constraints prohibiting following. Although “following conflicts” are defined in the literature, most solvers ignore them and it is non-trivial to incorporate them into certain MAPF algorithms. These combined factors mean that even though powerful, state-of-the-art solvers like LaCAM (Okumura 2023b, a, 2024) are effective for high-density MAPF problems, they are not directly applicable to BRaPs. Therefore, building upon these advanced frameworks, we develop a MAPF-based solution tailored to the specific complexities of the BRaP.
@@ 4.1 Configuration Space Search
The BRaP can be formulated as a graph search in configuration space. Starting from an initial grid configuration, available actions transform the state until reaching a goal configuration that satisfies desired criteria. Actions at each time step represent edges that transition between neighboring configuration states. Each state comprises time, assigned block set (vertices containing assigned blocks), empty vertex set, and completed block set (vertices containing completed assigned blocks).
An example graph of configuration states and actions is illustrated in Figure 4. According to the above definition, the root state in the graph is: time: 0, assigned blocks: {(2,5), (2,6)}, empty vertices: {(2,3), (2,4)}, completed blocks: $\emptyset$. The assigned block set in the goal state should be empty.
@@
The configuration space can be modified using the three actions described in Sec. 2.1. To ensure computational tractability and reduce the branching factor, we restrict the system to one action per time step. This restriction renders the wait action redundant and thus disallowed. Methods supporting multiple actions at each time step will be discussed in subsequent sections.
A configuration graph can be constructed from the defined states and actions, as shown in Figure 4. With the cost of the actions defined, graph search techniques like Dijkstra and A* algorithms can be applied to find the minimum cost path from the initial state to the goal state (where there is no longer any assigned blocks).
A* with an admissible heuristic function is guaranteed to be complete (always find a solution if there is) and optimal. We formulate an admissible heuristic by assuming assigned blocks move along least-blocking vertices, with each blocking block requiring only one move action for removal. The detailed heuristic formulation is presented in Appendix.
4.2 PDDL-based Configuration Space Search
The configuration space search can also be described using PDDL. We define five predicates: (emp ?u), (asb ?u), (blk ?u), (cmp ?u), and (goal ?u), indicating whether a vertex $u\in V$ contains an empty space, assigned block, unassigned block, completed block, or potential goal, respectively. The predicate (adjacent ?u ?v) denotes that vertices $u$ and $v$ are adjacent to each other, i.e., $(u,v)\in E$. Three actions are defined in the PDDL domain: (move_blk ?u ?v) and (move_asb ?u ?v) for moving unassigned and assigned blocks, respectively, and (complete ?u) for completion.
In the problem file, vertices serve as objects, denoted as node-0-0, node-0-1, etc., excluding vertices containing obstacles. The initial state specifies each node’s type as either (asb node-x-y), (blk node-x-y), or (emp node-x-y). All vertices except assigned blocks are marked with (cmp node-x-y), and valid terminal locations are marked with (goal node-x-y). The goal state requires all vertices to be marked as (cmp node-x-y), with total action cost as the optimization metric.
While this formulation allows simultaneous moves, the fast-downward solver used in Sec. 5 generates totally-ordered plans with single actions per timestep. Formulations that support multi-agent temporal planning would introduce prohibitively high branching factors for PDDL solvers, motivating the restriction to single-action plans.
@@ 4.3 Priority-based Configuration Space Search
In contrast to the single action per time step approach of Sec. 4.1, we now develop a formulation allowing one action per assigned block per time step, enabling parallel execution and reducing rearrangement makespan.
Algorithm 1 begins by assigning priorities based on block proximity to goals, as our experiments showed this produces better solutions. The priority for each assigned block $i\in\mathcal{I}$ is determined using the heuristic $h(s,i)$ described in Appendix, where smaller values indicate higher priorities.
The search iteratively creates rearrangement plans for individual assigned blocks while respecting plans of higher-priority blocks as constraints. For a prior plan $P$, the computeConstraint($P$) function generates constraints ensuring the preconditions of higher-priority actions are satisfied. For example, if move(i,t,u,v)(i,t,u,v) appears in $P$, a constraint in $\xi$ ensures block $i$ occupies $u$ at time $t$ while $v$ remains empty, and no other actions can modify these vertices at time $t$. The computeConstrainedPlan() function generates plans by pruning states violating $\xi$ and treating prior plans as external forces, ensuring parallel executability of the resulting plans.
@@
2Input: Grid $G(V,E)$, assigned block set $\mathcal{I}$, goal vertices $V_{i}(i\in\mathcal{I})$